BRICS Notes Series, 1997

July 7, 2003

This document is also available as PostScript and DVI.

References

[NS-97-1] Mogens Nielsen and Wolfgang Thomas, editors. Preliminary Proceedings of the Annual Conference of the European Association for Computer Science Logic, CSL ’97 (Aarhus, Denmark, August 23–29, 1997), August 1997. vi+432 pp.